In each part, compute f

, f

′′

, f

′′′ and then state the formula for nth derivative f

(n)


(a) f(x) = e

2x

(b) f(x) = sin3x

(c) f(x) = lnx

Draw the graph of first three derivatives of f(x).


1
Expert's answer
2021-10-24T18:06:15-0400

"(a)\ f(x) = e^{2x}\\
f'(x) = 2e^{2x}\\
f''(x) = 4e^{2x}\\
f'''(x) = 8e^{2x}\\

f^{(n)}(x) = 2^ne^{2x}"




graph of the first derivative




graph of the second derivative




graph of the third derivative




"(b)\ f(x) = sin3x\\
f'(x) = 3cos3x\\
f''(x) = -9sin3x\\
f'''(x) = -27cos3x\\
f^{(n)}(x) =3^n\sin\bigl(x+\tfrac{n\pi} 2\bigr)."




graph of the first derivative




graph of the second derivative




graph of the third derivative




"(c)\ f(x) = ln(x)\\
f'(x) = 1/x\\
f''(x) = -1/x²\\
f'''(x) = 2/x³\\
f^{(n)}(x) = \dfrac {({n - 1})!( {-1})^{n - 1} } {x^n}"




graph of the first derivative




graph of the second derivative




graph of the third derivative"."